OpenStudy (anonymous):

f(x)=9x-3 and g(x)=2x-4 find(f*g)(x)

OpenStudy (anonymous):

help

OpenStudy (jdoe0001):

\(\large {f(x)=9x-3 \qquad g(x)=2x-4\\ \quad \\ \begin{array}{cccl} f(x)=9x-3 && g(x)=2x-4\\ \uparrow &\times &\uparrow \\ f(x)&\times &g(x) \end{array}}\)

OpenStudy (jdoe0001):

(f*g)(x) = f(x) * g(x)

OpenStudy (jdoe0001):

or rather \(\large {f(x)=9x-3 \qquad g(x)=2x-4\\ \quad \\ \begin{array}{cccl} 9x-3 &\times& 2x-4\\ \uparrow & &\uparrow \\ f(x)&\times &g(x) \end{array}}\)